# Think Before You Diffuse: LLMs-Guided Physics-Aware Video Generation

### v1

@@ 4 Experiments
HQ-Phy Dataset. We curate a dataset of approximately 8,000 videos selected from the VIDGEN-1M dataset [46], which includes the video, caption, enhanced prompts, and extracted physical phenomena labels. We filter the dataset to exclude cartoon and other unrealistic scenes. To extract the enhanced prompts and detailed physical phenomena from this training dataset, we use DeepSeek-R1-32B [14] as a cost-effective alternative to GPT. Details are provided in Appendix.
Evaluation benchmarks: In our experiments, we utilize two datasets specifically designed to evaluate physical commonsense reasoning in text-to-video generation. VideoPhy2 [7] provides a collection of 590 human-verified captions that describe realistic interactions between various physical entities, including both solid and fluid objects. This dataset serves as a testbed for assessing whether generated videos reflect plausible real-world physics. Additionally, we incorporate PhyGenBench [37], a benchmark constructed to further probe the physical awareness of T2V models. It comprises 160 designed prompts spanning four fundamental domains of physics: mechanics, optics, thermal processes, and material properties.
@@
Metrics. To ensure a comprehensive evaluation of our generated videos, we adopt a multi-faceted assessment strategy combining model-based metrics and human judgment. Firstly, we adopt the evaluation metrics from the PhyGenBench benchmark [37]. Following PhyGenBench [37], we report key phenomena detection, physical order verification, and average scores to indicate overall quality using GPT-4o [1] and open-source models, i.e., CLIP [43], InternVideo2 [53], and LLaVA [32]. Second, we use the VideoCon-Physics evaluator from VideoPhy2[7], which reports two metrics: physical commonsense (PC) and semantic adherence (SA), each scored from 1 to 5. PC evaluates physical plausibility, while SA measures alignment with the input prompt. Following VideoPhy2[7], we set the overall score to 1 if both PC and SA are $\geq$ 4, and 0 otherwise. Finally, we conduct a gold-standard human evaluation via blind review, where three annotators independently score each video on semantic adherence and physical commonsense. This complements automatic metrics with qualitative insight.
Baselines. We apply DiffPhy on Wan2.1-14B [50], a DiT-based open-source T2V models, and compare it to eleven open or close-source T2V models, including Wan2.1-14B [50], CogVideoX-5B [63], VideoCrafter2 [10], HunyuanVideo-13B [27], Cosmos-Diffusion-7B [2], OpenAI Sora  [40], Luma Ray2 [35], LaVie [52], Vchitect [13], Pika [42], and Kling [26]. Although DiffPhy supports various backbones, we use Wan2.1-14B for illustration.
@@ 4.1 PhyGenBench Evaluation
Table 1 and Figure 4 present the performance of our model compared with a range of existing methods, including CogVideoX (2B/5B) [63], OpenSora [40], LaVie [52], Vichitect [13], Pika [42], Kling [26], and Wan 2.1-14B [50]. The results reported in Table 1 are sourced from the official PhyGenBench leaderboard [37], with the results for Wan 2.1-14B reproduced using its publicly released code and model [50]. Our DiffPhy not only outperforms Wan by a substantial margin across all metrics but also achieves the highest average score and the best performance in all of the dimensions, demonstrating the effectiveness of our proposed strategies.
Physical phenomena: We further validate the ability of our method to capture detailed physical phenomena. Table 2 and Figure 5 compare the performance of our model, the baseline Wan, and the second-ranked closed-source model Kling [26] on aspects including key phenomena (phenomena), physical order (order), and overall quality. The overall quality is indicated by the average results from GPT-4o [1] and open-sourced models (Open), i.e., CLIP [43], LLaVA [32], and InternVideo2 [53]. Our model achieves the highest scores in both key phenomena and physical order, with strong average performance across four physics-related categories: Mechanics, optics, thermal, and material properties. These results indicate that our model not only identifies key physical phenomena accurately but also generates videos that follow coherent physical progression.
@@ 4.2 Videophy bechmark
Table 4.1 and Figure 3 demonstrate that our method DiffPhy, significantly outperforms the existing text-to-video (T2V) approaches on the VideoPhy2 dataset. Compared to Wan, our model shows notable improvements in average performance across all subcategories, including sports and physical activities (Activity), object interactions (Interaction), and challenging physical scenarios (Hard). We evaluate our method on the VideoPhy2 benchmark, following the protocol outlined in [7]. To ensure a fair comparison, we report results both with and without CoT prompting. While our baseline already surpasses Wan, incorporating CoT further boosts the performance, achieving the highest average scores across all video categories. Figure 6 visualizes typical videos generated by our model with and without CoT reasoning, compared to the baseline Wan model. This demonstrates DiffPhy ’s ability to generate coherent and plausible videos. We also observe discrepancies between model-based and human-based ratings, which is also mentioned in [7]. This discrepancy arises because although model-based metrics capture some aspects of physical correctness, human evaluation remains the definitive standard for assessing physical plausibility. Results for the compared methods, including both model and human evaluations, are taken from the original VideoPhy2 leaderboard [7]. Our own human evaluation results are provided in Table 4.2.
Human evaluation. We conduct a human evaluation on two dimensions consistent with VideoPhy2 [7], including Semantic Alignment (SA) and Physical Commonsense (PC). Table 4.2 compares our model with the top two models on the Videophy2 human-annotated leaderboard, CogVideoX [17] and Wan 2.1-14B [51]. Our model outperforms Wan across all three evaluation dimensions, even without Chain-of-Thought (CoT). CoT further boosts performance across all dimensions.
@@
Table 5: Ablation results on VideoPhy2. $\mathcal{L}_{\text{com}}$, $\mathcal{L}_{\text{sem}}$,$\mathcal{L}_{\text{phen}}$ denote physical commonsense loss, semantic consistency loss, and physical phenomena loss, respectively; Iterative refers to an optional second-pass refinement by injecting the failure phenomena during inference.
@@
$\mathcal{L}_{\text{com}}$&$\mathcal{L}_{\text{sem}}$ $\mathcal{L}_{\text{phen}}$ Enhance Prompt IterativeActivityInteractionOverallActivityInteractionOverall
@@ 4.3 Ablations
We conduct ablation studies to evaluate the effectiveness of each design choices. Results are reported in Table 4.2. We have the following findings. First, incorporating MLLM-based loss functions yields substantial gains over the Wan [50] baseline on both automatic and human evaluation. Second, adding the attention injection to resolve failure facts with phenomena loss ($\mathcal{L}_{\text{phen}}$) further improves performance. Third, incorporating chain-of-thought (CoT) prompting enhances prompt clarity and reasoning, resulting in additional improvement. Visual results on the effects of CoT are reported in Figure 6. During inference, our DiffPhy framework supports an optional two-pass inference by correcting failure cases from previous pass via attention injection, which yields additional performance improvements. Because it introduces runtime overhead, for fair comparison, we only report results using one-pass inference for the main experiments.
@@ Appendix B Implementation Details
[…]

### v2

@@ 4 Experiments
HQ-Phy Dataset. We curate a dataset of approximately 8,000 videos selected from the VIDGEN-1M dataset tan2024vdgen-1m , which includes the video, caption, enhanced prompts, and extracted physical phenomena labels. We filter the dataset to exclude cartoon and other unrealistic scenes. To extract the enhanced prompts and detailed physical phenomena from this training dataset, we use DeepSeek-R1-32B guo2025deepseek as a cost-effective alternative to GPT. Details are provided in Appendix.
Evaluation benchmarks: In our experiments, we utilize two datasets specifically designed to evaluate physical commonsense reasoning in text-to-video generation. VideoPhy2 bansal2025videophy provides a collection of 590 human-verified captions that describe realistic interactions between various physical entities, including both solid and fluid objects. This dataset serves as a testbed for assessing whether generated videos reflect plausible real-world physics. Additionally, we incorporate PhyGenBench meng2024towards , a benchmark constructed to further probe the physical awareness of T2V models. It comprises 160 designed prompts spanning four fundamental domains of physics: mechanics, optics, thermal processes, and material properties.
@@
Metrics. To ensure a comprehensive evaluation of our generated videos, we adopt a multi-faceted assessment strategy combining model-based metrics and human judgment. Firstly, we adopt the evaluation metrics from the PhyGenBench benchmark meng2024towards . Following PhyGenBench meng2024towards , we report key phenomena detection, physical order verification, and average scores to indicate overall quality using GPT-4o achiam2023gpt4 and open-source models, i.e., CLIP radford2021learning , InternVideo2 wang2024internvideo2 , and LLaVA lin2023videoLLaVA . Second, we use the VideoCon-Physics evaluator from VideoPhy2bansal2025videophy , which reports two metrics: physical commonsense (PC) and semantic adherence (SA), each scored from 1 to 5. PC evaluates physical plausibility, while SA measures alignment with the input prompt. Following VideoPhy2bansal2025videophy , we set the overall score to 1 if both PC and SA are $\geq$ 4, and 0 otherwise. Finally, we conduct a gold-standard human evaluation via blind review, where three annotators independently score each video on semantic adherence and physical commonsense. This complements automatic metrics with qualitative insight.
Baselines. We apply DiffPhy on Wan2.1-14B wan14b , a DiT-based open-source T2V models, and compare it to eleven open or close-source T2V models, including Wan2.1-14B wan14b , CogVideoX-5B yang2024cogvideox , VideoCrafter2 chen2024videocrafter2 , HunyuanVideo-13B kong2024hunyuanvideo , Cosmos-Diffusion-7B agarwal2025cosmos , OpenAI Sora  openai2024sora , Luma Ray2 lumalabsLumaRay2 , LaVie wang2023laviehighqualityvideogeneration , Vchitect fan2025vchitect , Pika pika2025 , and Kling klingai . Although DiffPhy supports various backbones, we use Wan2.1-14B for illustration.
@@ 4.1 PhyGenBench Evaluation
Table 1 and Figure 4 present the performance of our model compared with a range of existing methods, including CogVideoX (2B/5B) yang2024cogvideox , OpenSora openai2024sora , LaVie wang2023laviehighqualityvideogeneration , Vichitect fan2025vchitect , Pika pika2025 , Kling klingai , and Wan 2.1-14B wan14b . The results reported in Table 1 are sourced from the official PhyGenBench leaderboard meng2024towards , with the results for Wan 2.1-14B reproduced using its publicly released code and model wan14b . Our DiffPhy not only outperforms Wan by a substantial margin across all metrics but also achieves the highest average score and the best performance in all of the dimensions, demonstrating the effectiveness of our proposed strategies.
Physical phenomena: We further validate the ability of our method to capture detailed physical phenomena. Table 2 and Figure 5 compare the performance of our model, the baseline Wan, and the second-ranked closed-source model Kling klingai on aspects including key phenomena (phenomena), physical order (order), and overall quality. The overall quality is indicated by the average results from GPT-4o achiam2023gpt4 and open-sourced models (Open), i.e., CLIP radford2021learning , LLaVA lin2023videoLLaVA , and InternVideo2 wang2024internvideo2 . Our model achieves the highest scores in both key phenomena and physical order, with strong average performance across four physics-related categories: Mechanics, optics, thermal, and material properties. These results indicate that our model not only identifies key physical phenomena accurately but also generates videos that follow coherent physical progression.
@@ 4.2 Videophy bechmark
Table 4.1 and Figure 3 demonstrate that our method DiffPhy, significantly outperforms the existing text-to-video (T2V) approaches on the VideoPhy2 dataset. Compared to Wan, our model shows notable improvements in average performance across all subcategories, including sports and physical activities (Activity), object interactions (Interaction), and challenging physical scenarios (Hard). We evaluate our method on the VideoPhy2 benchmark, following the protocol outlined in bansal2025videophy . To ensure a fair comparison, we report results both with and without CoT prompting. While our baseline already surpasses Wan, incorporating CoT further boosts the performance, achieving the highest average scores across all video categories. Figure 6 visualizes typical videos generated by our model with and without CoT reasoning, compared to the baseline Wan model. This demonstrates DiffPhy ’s ability to generate coherent and plausible videos. We also observe discrepancies between model-based and human-based ratings, which is also mentioned in bansal2025videophy . This discrepancy arises because although model-based metrics capture some aspects of physical correctness, human evaluation remains the definitive standard for assessing physical plausibility. Results for the compared methods, including both model and human evaluations, are taken from the original VideoPhy2 leaderboard bansal2025videophy . Our own human evaluation results are provided in Table 4.2.
Human evaluation. We conduct a human evaluation on two dimensions consistent with VideoPhy2 bansal2025videophy , including Semantic Alignment (SA) and Physical Commonsense (PC). Table 4.2 compares our model with the top two models on the Videophy2 human-annotated leaderboard, CogVideoX hong2022cogvideo and Wan 2.1-14B wang2025wan . Our model outperforms Wan across all three evaluation dimensions, even without Chain-of-Thought (CoT). CoT further boosts performance across all dimensions.
@@
Table 5: Ablation results on VideoPhy2. $\mathcal{L}_{\text{com}}$, $\mathcal{L}_{\text{sem}}$$\mathcal{L}_{\text{phen}}$ denote physical commonsense loss, semantic consistency loss, and physical phenomena loss, respectively; Iterative refers to an optional second-pass refinement by injecting the failure phenomena during inference.
@@
$\mathcal{L}_{\text{com}}$$\mathcal{L}_{\text{sem}}$ $\mathcal{L}_{\text{phen}}$ Enhance Prompt IterativeActivityInteractionOverallActivityInteractionOverall
@@ 4.3 Ablations
We conduct ablation studies to evaluate the effectiveness of each design choices. Results are reported in Table 4.2. We have the following findings. First, incorporating MLLM-based loss functions yields substantial gains over the Wan wan14b baseline on both automatic and human evaluation. Second, adding the attention injection to resolve failure facts with phenomena loss ($\mathcal{L}_{\text{phen}}$) further improves performance. Third, incorporating chain-of-thought (CoT) prompting enhances prompt clarity and reasoning, resulting in additional improvement. Visual results on the effects of CoT are reported in Figure 6. During inference, our DiffPhy framework supports an optional two-pass inference by correcting failure cases from previous pass via attention injection, which yields additional performance improvements. Because it introduces runtime overhead, for fair comparison, we only report results using one-pass inference for the main experiments.
@@ Appendix B Implementation Details
The multimodal LLM is used to guide the model in generating physically plausible videos during training. We adopt the multimodal LLM evaluator VideoCon-Physics [7] for its lightweight, open-source design. It is built on VideoCon [5], with CLIP [43] as the visual encoder and LLaMA-7B [47] as the language backbone. For MLLM Output Processing, due to the inherent noise and uncertainty in LLM-generated responses, we apply a post-processing strategy to stabilize supervision by filtering out invalid word tokens. Specifically, we use the pretrained multimodal LLM evaluator to predict discrete score distributions: integers from 1 to 5 for semantic consistency and physical commonsense, and labels $0:\text{unfollow''},1:\text{follow''},2:\text{``undetermined''}$} for each physical phenomenon. Since the multimodal LLM outputs textual tokens rather than strict integers, some predictions may be noisy or invalid. We therefore filter out invalid tokens and retain only those that correspond to the valid integer labels.